How do i find an equation of each line with the given slope that passes through the given point and Write the equation in the form. Ax+ By= C
m= - 2/9; (2,5)

Found 2 solutions by mananth, solver91311:
Answer by mananth(16949)

m= - 2/ 9

Plug value of the slope and point (2,5) in y=mx+b
5.00 = - 2/9*2 +b
b= 5 + 4/9

b= 49/ 9
So the equation will be
Y = - 2/9 x + 49/9
Multiply equation by 9
9y=-2x+49
2x+9y=49
m.ananth@hotmail.ca

Answer by solver91311(24713)



Use the point-slope form of an equation of a line:



where are the coordinates of the given point and is the given/calculated slope.

Then do the arithmetic and rearrange to get to the form.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism